# Only Send What You Need: Learning to Communicate Efficiently in Federated Multilingual Machine Translation

### v1

@@ 4.3 FL Training and MAML Module
We build our FL experiments using the Flower framework Beutel et al. (2020) for training and evaluation. For centralized experiments, we train models for 50 epochs and discuss the effect of pre-trained knowledge for NMT. For every FL experiment, we train each method for 25 communication rounds (epochs) and initialize the clients using a pre-trained M2M-100 model from Hugging Face’s transformers library Wolf et al. (2019). As a reference, we also conduct FL experiments by initializing the clients’ model with random weights; the corresponding results can be found in Appendix G.
@@
For our MAML module, we use a multi-layer perceptron (MLP) network with one hidden layer containing 100 neurons as the default setting. The ablation study in Section 5.3 presents the results of MetaSend considering different numbers of neurons in the MAML module. To randomly sample a small portion of the validation set for the MAML update, we use 16 batches (i.e., ). In Section 5.3, we also present an ablation study on to see the effect of MAML optimization for NMT quality. See Appendix B for detailed hyperparameters and compute settings.
4.4 Baselines
@@
We use several competitive baseline approaches and parameter selection strategies to evaluate federated NMT. PMFL Weller et al. (2022) is the basic FL framework that uses a pre-trained model for federated NMT without any decision-making mechanism. and are the recent methods from Passban et al. (2022) that select which tensors to send by comparing the norm difference between the previous and current client models. Their thresholding mechanism sorts based on the norm difference and either send the top 50% () or bottom 50% () of tensors during the aggregation process. We also include the results from a random configuration, RandSend, which randomly sends 50% of the tensors during FL aggregation.
5 Experimental Results
5.1 Translation Performance Evaluation
@@
Table 1 presents the SacreBLEU results of the translation task for both datasets. In the first section, we observe that the centralized method outperforms PMFL methods, as we would expect; on the other hand, it compromises data privacy by not preserving individual client data confidentiality. Further, the performance decrease of PMFL from IID to non-IID FL training reveals the challenges in the practical NMT scenario of clients having only single language directions.
@@
Among all the parameter selection methods, achieves the highest scores in both sacreBLEU and COMET metrics (refer to Appendix C). It demonstrates comparable translation quality to PMFL, indicating its ability to preserve communication resources without compromising translation quality. Translation examples generated by our method and the baseline are provided in Appendix E, where it is evident that our method shows better alignment with the ground truth regarding sentiment and accurate word usage. Additional results presented in Appendix F show the consistent superiority of our method over the baselines, not only in scenarios with ample training data but also in situations where clients have limited data resources.
5.2 Communication Efficiency Evaluation
@@
In Figure 4, we present the average sacreBLEU score and tensor savings for each method across 25 communication rounds on the MTNT (a and b) and UNMT (c and d) datasets. The RandSend and DP methods can save around 50% of tensors during FL communication due to their designs. By sending the tensors based on a specific learned threshold, MetaSend methods obtain substantial improvements in translation quality compared to DP methods, while also obtaining varying degrees of tensor savings. Among our two MetaSend methods, MetaSend outperforms in translation quality, indicating that sending the majority of tensors for update ensures significant performance improvement. On the other hand, MetaSend demonstrates higher tensor savings, with an average of 10.3% more tensors saved compared to DP methods.
In addition to evaluating tensor savings, Table 2 provides the specific training time for each method. During local iterations, all methods require a similar amount of time to process the entire local training dataset and update all parameters in the model. Although PMFL does not spend time on the parameter selection process, it consumes the most time in client communication and aggregation due to the necessity of transmitting and aggregating every tensor in the model. Both DP and MetaSend require computation time for calculating deviations between a client’s current tensors and its tensors from the previous round. However, DP carries out its operation by selecting either the top or bottom 50% of tensors, which occurs after all deviations have been calculated and sorted. In contrast, MetaSend immediately decides whether to send a tensor based on the learned threshold after calculating a single deviation and without any sorting calculation. As a result, MetaSend requires less time for parameter selection compared to DP. Finally, the time required for sending and aggregation depends on the tensor savings of each scheme.
@@
Both MetaSend and MetaSend involve additional computation for the MAML module, and they spend a similar amount of time on this module as it is independent of the operator. The breakdown of the time spent on the MAML module for MetaSend methods is provided in Table 3. We observe that the most time-consuming configuration is the meta-evaluation, which requires forward-passing a few batches to the global model and obtaining the validation loss. Therefore, we conduct an ablation study in Section 5.3 to examine the impact of the number of sampled batches. In sum, our proposed MetaSend significantly enhances translation quality while achieving greater resource savings compared to both PMFL and DP methods.
@@
5.3 Ablation Studies
@@
Effectiveness of learned threshold. To isolate the impact of the sending threshold, we compare MetaSend with different thresholds, including our learned threshold, a fixed threshold (), and a random threshold selected from 0 to 1. The red arrows in Figure 5 show the improvements in MetaSend when using different thresholds within a single operator ( or ). Compared with other thresholds, our learned threshold significantly increases both translation quality and the amount of tensor savings. By comparing Figures 5 and 4, MetaSend with a fixed threshold sometimes outperforms DP methods, suggesting that sending based on the deviation distribution should be considered instead of sending approximately half of the model’s tensors. See Appendix D for additional details and a comprehensive discussion on our learned threshold.
@@
Parameters used in MAML module. To explore the impact of the number of neurons on performance, we keep the learning parameters consistent while varying the number of neurons in the hidden layer of the MAML module. In Figure 6, we see that using more neurons in the MAML module generally leads to improved results in terms of sacreBLEU score and tensor savings. The performance gain from using more neurons is intuitive since it provides additional degrees of freedom for learning optimization. However, it is important to note that using more neurons also incurs higher resource requirements during system construction.
@@
Meta evaluation for MAML module. To examine the influence of the number of batches used to optimize our MAML module, Figure 7 shows the performance of our method with different numbers of batches used for optimization. We see that increasing the number of samples used for MAML optimization generally results in improved translation quality and efficiency. Naturally, using more batches will increase the exact time spent on our MAML module. However, the time taken by clients to send parameters for aggregation may be more critical than this optimization time, as the optimization process is performed only once in each round.
Additional experiments. Other results on limited training data and the absence of initialized knowledge, as well as detailed discussions on our learned threshold and ablation studies with other datasets and data distributions, are available in Appendices D–F.
6 Conclusion
We conducted the first analysis on the efficiency of federated multilingual NMT. To address the practical challenges that arise in this setup, we proposed MetaSend, which selects tensors for transmission that are most critical to the NMT. By adaptively learning the sending threshold in each FL round based on meta-learning, we saw that MetaSend not only improves communication efficiency, but also effectively captures the NMT threshold for sending. Extensive experiments on two datasets showed that MetaSend outperforms existing baselines regarding machine translation quality and significantly reduces communication costs in FL, confirming its advantage in practical federated NMT settings.
Appendix A Details for Datasets and Evaluation Metrics
This paper considers two widely used NMT datasets: MTNT Michel and Neubig (2018) and UN Corpus Ziemski et al. (2016). The Machine Translation of Noisy (MTNT) dataset Michel and Neubig (2018) was gathered from user comments on Reddit discussion threads. The dataset contains two language directions: English to French (En Fr) and English to Japanese (En Ja). The dataset contains 5,605 instances in each direction for training and approximately 1k each for validation and test set. The UN Corpus Ziemski et al. (2016) consists of manually translated UN documents over the years 1990 to 2014, and we consider three official UN language directions: English to French (En Fr), Arabic to Spanish (Ar Es), and Russian to Chinese (Ru Zh). The dataset contains 80k instances in each direction for training and approximately 10k each for validation and test set.
The evaluation metric sacreBLEU is widely used in the machine translation community and is built upon BLEU Papineni et al. (2002). Our study uses the standard sacreBLEU settings, including nrefs:1, mixed case, eff:no, tok:13a, smooth:exp, and version 2.0.0. For, Japanese (Ja) and Chinese (Zh), we use their respective tokenizers to ensure accurate evaluation. We utilize the default COMET model as suggested by the authors, which employs a reference-based regression approach and is developed based on XLM-R. This model has been trained on direct assessments from WMT17 to WMT20 and assigns scores ranging from 0 to 1, where 1 indicates a perfect translation. The COMET metric has been found to exhibit segment-level correlation with human evaluations and has demonstrated its potential to distinguish between high-performing systems more effectively.
Appendix B Hyperparameters and Compute Settings
For all the experiments, the MT engine (M2M-100 model) is optimized using the Adam optimizer. We search the learning rates from [1e-2, 5e-3, 1e-3, 5e-4, 1e-4] and select 5e-3 as the optimal learning rate for the MT engine. The batch size for the MT engine is set to 2 for both client training and updating the MAML module. The MAML module is optimized via Adam optimizer with a learning rate 1e-3, which is also searched from [1e-2, 5e-3, 1e-3, 5e-4, 1e-4].
We run all experiments on a 3-GPU cluster of Tesla V100 GPUs, with each GPU having 32GB of memory. Centralized experiments will be conducted on one of our 3 Tesla V100 GPUs, while FL experiments will utilize GPUs, where is the total number of clients. Each centralized experiment ran for approximately four hours and two days for the NTMT and UNMT datasets, respectively, when run on a single GPU. For FL experiments on the NTMT dataset, each simulation was completed in about 2 hours by distributing clients’ data on 2 GPUs. For FL experiments on the UNMT dataset, each simulation was run for approximately 14 hours by distributing clients’ data on 3 GPUs.
Appendix C Translation Quality Evaluated by COMET
This section provides supplementary results to Section 5.1. Several research works have mentioned that the traditional overlap-based evaluation metrics do not correlate well with human evaluation Sinha et al. (2020); Zhang et al. (2020); Sellam et al. (2020); Hsu et al. (2022, 2021). Therefore, we also evaluated each method using a novel MT evaluation metric, COMET Rei et al. (2020), demonstrating that the evaluated quality is better aligned with human judgement. Table 4 shows the translation quality evaluated using COMET for each method. Consistent with the sacreBLEU results in Table 1, MetaSend demonstrates the highest average COMET score across two datasets, while MetaSend achieves comparable quality.
| Training | Method | MTNT | UNMT | |||||
| En Fr | En Ja | Avg | En Fr | Ar Es | Ru Zh | Avg | ||
| IID FL | PMFL | 0.758 | 0.734 | 0.746 | 0.853 | 0.839 | 0.830 | 0.841 |
| Non-IID FL | PMFL | 0.666 | 0.659 | 0.663 | 0.737 | 0.742 | 0.709 | 0.729 |
| IID FL | RandSend | 0.726 | 0.715 | 0.721 | 0.811 | 0.773 | 0.811 | 0.798 |
| 0.729 | 0.721 | 0.725 | 0.825 | 0.801 | 0.823 | 0.816 | ||
| 0.737 | 0.730 | 0.734 | 0.830 | 0.805 | 0.829 | 0.821 | ||
| MetaSend (ours) | 0.755 | 0.733 | 0.744 | 0.846 | 0.829 | 0.833 | 0.836 | |
| MetaSend (ours) | 0.755 | 0.736 | 0.746 | 0.849 | 0.827 | 0.831 | 0.836 | |
| 0.651 | 0.650 | 0.651 | 0.709 | 0.720 | 0.684 | 0.704 | ||
| 0.659 | 0.648 | 0.6534 | 0.713 | 0.729 | 0.692 | 0.711 | ||
| MetaSend (ours) | 0.668 | 0.653 | 0.661 | 0.744 | 0.739 | 0.713 | 0.732 | |
| MetaSend (ours) | 0.666 | 0.659 | 0.663 | 0.749 | 0.744 | 0.721 | 0.738 | |
Appendix D Learned Threshold of MetaSend
This section provides supplementary results to Section 5.3, and Figure 8 shows the learned threshold for our method. The decreasing value of for indicates its focus on sending tensors with even smaller deviations during communication. As the model converges and becomes stable, the deviations decrease, resulting in generating a smaller threshold for a milder deviation distribution.
Appendix E Case Study
This section provides some translation examples to Section 5.1. Figure 9 illustrates that our method generates translations that closely align with the ground truth by utilizing similar words. In contrast, the baseline method produces translations with redundant tokens, leading to potential confusion within the sentence. Figure 10 shows that our method employs the same words as the ground truth, conveying a neutral sentiment. In comparison, the baseline method generates similar words but with a negative sentiment.
Figure 11 shows the challenge of training FL algorithm under Non-IID data distribution. We can see that our method, trained under the IID setting, achieves a good translation compared to the ground truth by round 9. However, when trained under the Non-IID setting, our method requires 17 communication rounds to achieve comparable results. The result obtained at round 9 for our method trained under the Non-IID setting shows that the model has not yet converged, resulting in chaotic translations with mixed languages.
Appendix F Insufficient Training Samples for Clients
To mirror the limited data scenario of each client in practical FL, we performed experiments by randomly sampling a small portion of data from MTNT and UNMT datasets as training data. Specifically, we randomly selected 20% of samples from each dataset, resulting in 1k and 10k training samples in each language direction for MTNT and UNMT datasets, respectively, while keeping the validation and test sets at the same size. All other hyperparameters, such as batch size for the NMT engine or MAML optimization, neurons in the MAML module, and learning rate, remain the same as in Section 5.
Tables 5 and 6 present the translation quality of each method when trained with limited data. It is evident that our MetaSend methods consistently outperform other baselines and sometimes deliver comparable or even slightly better performance compared to PMFL. shows overall superiority in the UNMT dataset, whereas the results in the MTNT dataset are more diverse due to the presence of noise in the dataset.
| Training | Method | MTNT | UNMT | |||||
| En Fr | En Ja | Avg | En Fr | Ar Es | Ru Zh | Avg | ||
| IID FL | PMFL | 27.3 | 11.5 | 19.4 | 33.6 | 33.2 | 34.4 | 33.7 |
| Non-IID FL | PMFL | 16.6 | 8.8 | 12.7 | 19.4 | 19.6 | 17.7 | 18.9 |
| IID FL | RandSend | 26.1 | 10.4 | 18.3 | 31.5 | 31.3 | 32.1 | 31.6 |
| 27.3 | 10.9 | 19.1 | 32.6 | 32.7 | 32.3 | 32.5 | ||
| 26.8 | 11.0 | 18.9 | 33.0 | 32.9 | 32.8 | 32.9 | ||
| MetaSend (ours) | 28.0 | 11.8 | 19.9 | 34.3 | 33.0 | 34.1 | 33.8 | |
| MetaSend (ours) | 27.3 | 12.0 | 19.7 | 34.2 | 33.2 | 34.1 | 33.8 | |
| Non-IID FL | RandSend | 14.9 | 7.4 | 11.1 | 18.2 | 17.4 | 16.4 | 17.3 |
| 15.3 | 7.9 | 11.6 | 18.4 | 19.5 | 16.1 | 18.0 | ||
| 15.0 | 7.3 | 11.2 | 18.9 | 19.0 | 16.4 | 18.1 | ||
| MetaSend (ours) | 15.8 | 7.9 | 11.9 | 19.5 | 19.5 | 17.5 | 18.8 | |
| MetaSend (ours) | 16.0 | 7.8 | 11.9 | 19.4 | 19.7 | 17.7 | 18.9 | |
| Training | Method | MTNT | UNMT | |||||
| En Fr | En Ja | Avg | En Fr | Ar Es | Ru Zh | Avg | ||
| IID FL | PMFL | 0.687 | 0.683 | 0.685 | 0.750 | 0.736 | 0.745 | 0.744 |
| Non-IID FL | PMFL | 0.640 | 0.661 | 0.651 | 0.677 | 0.664 | 0.671 | 0.671 |
| IID FL | RandSend | 0.666 | 0.655 | 0.661 | 0.733 | 0.719 | 0.727 | 0.726 |
| 0.701 | 0.678 | 0.690 | 0.746 | 0.727 | 0.739 | 0.737 | ||
| 0.691 | 0.680 | 0.686 | 0.741 | 0.729 | 0.733 | 0.734 | ||
| MetaSend (ours) | 0.711 | 0.680 | 0.696 | 0.752 | 0.736 | 0.743 | 0.744 | |
| MetaSend (ours) | 0.687 | 0.684 | 0.685 | 0.750 | 0.736 | 0.747 | 0.743 | |
| 0.633 | 0.654 | 0.644 | 0.673 | 0.649 | 0.659 | 0.660 | ||
| 0.631 | 0.659 | 0.645 | 0.680 | 0.653 | 0.678 | 0.670 | ||
| MetaSend (ours) | 0.640 | 0.670 | 0.655 | 0.659 | 0.670 | 0.711 | 0.680 | |
| MetaSend (ours) | 0.638 | 0.670 | 0.654 | 0.685 | 0.688 | 0.685 | 0.686 | |
Appendix G FL Experiments Without Using Pre-trained Knowledge
Tables 8 and 9 show the results of each method without utilizing pre-trained knowledge for the M2M-100 model in each client. Specifically, each method was trained from scratch, utilizing randomly initialized weights. Although the results of do not exhibit complete superiority over as discussed in Section 5.1, our MetaSend methods display higher average scores overall in most cases compared with baseline approaches. However, the overall results are relatively poor and unstable compared to the results shown in Tables 1 and 4 when pre-training is utilized since large neural MT systems generally require a substantial amount of data to perform effectively.

### v2

@@ IV-C FL Training and MAML Module
We build our FL experiments using the Flower framework [46] for training and evaluation. For centralized experiments, we train models for 50 epochs and discuss the effect of pre-trained knowledge for NMT. For every FL experiment, we train each method for 25 communication rounds (epochs) and initialize the clients using a pre-trained M2M-100 model from Hugging Face’s transformers library [47]. As a reference, we also conduct FL experiments by initializing the clients’ model with random weights.
@@
For our MAML module, we use a multi-layer perceptron (MLP) network with one hidden layer containing 100 neurons as the default setting [48]. The ablation study in Section V-C2 presents the results of MetaSend considering different numbers of neurons in the MAML module. To randomly sample a small portion of the validation set for the MAML update, we use 16 batches (i.e., ). In Section V-C3, we also present an ablation study on to see the effect of MAML optimization for NMT quality.
IV-D Baselines
@@
We use several competitive baseline approaches and parameter selection strategies compare with MetaSend in federated NMT. PMFL [10] is the basic FL framework that uses a pre-trained model for federated NMT without any decision-making mechanism. and are recent methods from [14] that select which tensors to send by comparing the norm difference between the previous and current client models. Their thresholding mechanism first sorts the parameters by norm differences and either sends the top 50% () or the bottom 50% () of parameters for aggregation. We also include the results from a random configuration, RandSend, which randomly sends 50% of the tensors during FL aggregation. In addition to the baselines designed for federated multilingual NMT,we compare our method to top-k sparsification [25] and Adaptive Parameter Freezing (APF) [29], two recent communication-efficient FL methods intended for more generic ML tasks. Top-k sparsification sorts the model parameters based on their absolute values and sends only the largest k parameters for aggregation. APF employs an observation window consisting of a certain number of consecutive model updates from recent past rounds to compute perturbations and determine which parameters need to be communicated.
V Experimental Results
V-A Translation Performance Evaluation
| Training | Method | MTNT | UNMT | |||||
| En Fr | En Ja | Avg | En Fr | Ar Es | Ru Zh | Avg | ||
| IID FL | PMFL | 0.758 | 0.734 | 0.746 | 0.853 | 0.839 | 0.830 | 0.841 |
| Non-IID FL | PMFL | 0.666 | 0.659 | 0.663 | 0.737 | 0.742 | 0.709 | 0.729 |
| IID FL | RandSend | 0.726 | 0.715 | 0.721 | 0.811 | 0.773 | 0.811 | 0.798 |
| DPg | 0.729 | 0.721 | 0.725 | 0.825 | 0.801 | 0.823 | 0.816 | |
| DPl | 0.737 | 0.730 | 0.734 | 0.830 | 0.805 | 0.829 | 0.821 | |
| MetaSendg (ours) | 0.755 | 0.733 | 0.744 | 0.846 | 0.829 | 0.833 | 0.836 | |
| MetaSendl (ours) | 0.755 | 0.736 | 0.746 | 0.849 | 0.827 | 0.831 | 0.836 | |
| DPg | 0.651 | 0.650 | 0.651 | 0.709 | 0.720 | 0.684 | 0.704 | |
| DPl | 0.659 | 0.648 | 0.653 | 0.713 | 0.729 | 0.692 | 0.711 | |
| MetaSendg (ours) | 0.668 | 0.653 | 0.661 | 0.744 | 0.739 | 0.713 | 0.732 | |
| MetaSendl (ours) | 0.666 | 0.659 | 0.663 | 0.749 | 0.744 | 0.721 | 0.738 | |
@@ V-A Translation Performance Evaluation
Table I and II present the SacreBLEU and COMET results of the translation task for both MTNT and UNMT datasets. In the first section of the tables, we observe that the centralized method outperforms PMFL methods, as we would expect; on the other hand, it compromises data privacy by not preserving individual client data confidentiality. Further, the performance decrease of every method from IID to non-IID FL training reveals the challenges in the practical NMT scenario of clients having only single language directions. Figure 4 shows the challenge of training FL algorithm under Non-IID data distribution. We can see that our method, trained under the IID setting, achieves a good translation compared to the ground truth by round 9. However, when trained under the Non-IID setting, our method requires 17 communication rounds to achieve comparable results. The result obtained at round 9 for our method trained under the Non-IID setting shows that the model has not yet converged, resulting in chaotic translations with mixed languages.
@@
Among all the parameter selection methods, achieves the highest scores in both sacreBLEU and COMET metrics. It demonstrates comparable translation quality to PMFL, indicating its ability to preserve communication resources without compromising translation quality. Translation examples generated by our method and the baseline are provided in Figure 5, where it is evident that our method shows better alignment with the ground truth regarding sentiment and accurate word usage. Figure 5(a) illustrates that our method generates translations that closely align with the ground truth by utilizing similar words. In contrast, the baseline method produces translations with redundant tokens, leading to potential confusion within the sentence. Figure 5(b) shows that our method employs the same words as the ground truth, conveying a neutral sentiment. In comparison, the baseline method generates similar words but with a negative sentiment.
In addition to the baselines DPl and DPg that are explicitly designed for federated NMT scenarios, we also evaluate the performance of top- sparsification and APF in this language domain. Table III presents the average translation results along with the parameter savings obtained by our method and these baselines across both datasets and training settings. The value of is set according to 50% of the largest model parameters for top- compression, and the historical window size for APF is set to 4. The results demonstrate that our MetaSendl achieves the highest average translation quality, while our MetaSendg consistently outperforms top- compresion and APF in both translation performance and communication efficiency. This highlights the effectiveness of our MAML-based adaptive thresholding technique, which is designed to address the nuances of federated multilingual NMT. Moreover, while these parameter-wise methods focus on communication efficiency, they fail to account for the unique deviation patterns in NMT, which can result in sending either too many or too few parameters, missing opportunities for improving either communication efficiency or translation performance without noticeably affecting the other metric. The thresholds reflected by these approaches are based solely on parameter magnitude or past parameters and are not optimized with consideration for the model’s performance in federated NMT, as our method achieves. These differences highlight the advantages of our approach in balancing communication efficiency and maintaining model performance for multilingual NMT tasks.
| Training | Method | MTNT | UNMT | ||
| Translation Performance | Tensor Saving | Translation Performance | Tensor Saving | ||
| IID FL | top-k compression | 23.0 | 50.0% | 38.2 | 50.1% |
| APF | 23.9 | 58.7% | 41.0 | 52.7% | |
| MetaSendg (ours) | 24.4 | 61.5% | 41.0 | 59.2% | |
| MetaSendl (ours) | 25.0 | 41.9% | 41.1 | 44.5% | |
| Non-IID FL | top-k compression | 15.7 | 49.9% | 28.2 | 49.9% |
| APF | 16.4 | 49.3% | 29.4 | 41.2% | |
| MetaSendg (ours) | 16.8 | 62.3% | 30.3 | 58.3% | |
| MetaSendl (ours) | 16.9 | 42.7% | 31.3 | 43.9% | |
V-B Communication Efficiency Evaluation
@@
In addition to translation quality, recall that one of our key objectives is improving efficiency during FL communication. In Figure 6, we present the average sacreBLEU score and tensor savings for each method across 25 communication rounds on the MTNT (a and b) and UNMT (c and d) datasets. The RandSend and DP methods can save around 50% of tensors during FL communication due to their designs. By sending the tensors based on a specific learned threshold, MetaSend methods obtain substantial improvements in translation quality compared to DP methods, while also obtaining varying degrees of tensor savings. Among our two MetaSend methods, MetaSendl outperforms in translation quality, indicating that sending the majority of tensors for update ensures significant performance improvement. On the other hand, MetaSendg demonstrates higher tensor savings, with an average of 10.3% more tensors saved compared to DP methods.
Figure 7 visualizes the translation quality and parameter savings obtained over federated training rounds for DP and our method, under the non-IID FL scenario. This offers a more detailed perspective on performance and communication efficiency throughout the training process. We can see that MetaSendg consistently provides efficiency benefits over the DP method across all rounds. Recall from Figure 1 that tensor distributions over FL rounds tend to cluster predominantly in low-variation regions, with some tensors falling into highly active areas. MetaSendg, which selects parameters greater than the dynamically learned threshold, often transmits fewer parameters, resulting in higher tensor savings across rounds. In contrast, MetaSendl, which selects parameters below the threshold, typically transmits more parameters, leading to lower tensor savings. These differences highlight how the dynamic thresholds in MetaSend adapt to varying tensor deviations, determining crucial parameters for communication based on the machine translation performance.
In addition to evaluating tensor savings, Table IV reports training times for each method. All methods require similar time for local iterations, processing the full dataset and updating all parameters. While PMFL skips parameter selection, it spends the most time on communication and aggregation as it transmits and aggregates all tensors. Both DP and MetaSend compute deviations between a client’s current and previous round tensors. While DP sorts all deviations and selects the top or bottom 50% of tensors, MetaSend uses a learned threshold to decide transmissions without sorting, resulting in faster parameter selection. Additionally, MetaSendg, which transmits the fewest parameters, requires the least time for sending and aggregation.
@@
Both MetaSendl and MetaSendg involve additional computation for the MAML module, and they spend a similar amount of time on this module as it is independent of the operator. Table V provides a breakdown of time spent on the MAML module for MetaSend. We observe that meta-evaluation is the most time-consuming, involving forward passes to the global model and computing the validation loss. Therefore, we conduct an ablation study in Section V-C3 to examine the impact of the number of sampled batches. In sum, our proposed MetaSend significantly enhances translation quality while achieving greater resource savings compared to both PMFL and DP methods.
@@
V-C Varying System Parameters and Configuration
V-C1 Effectiveness of learned threshold
@@
To isolate the impact of the sending threshold, we compare MetaSend with different thresholds, including our learned threshold, a fixed threshold (), and a random threshold selected from 0 to 1. The red arrows in Figure 8 show the improvements in MetaSend when using different thresholds within a single operator ( or ). Compared with other thresholds, our learned threshold improves both translation quality and tensor savings. By comparing Figures 8 and 6, MetaSend with a fixed threshold sometimes outperforms DP methods, showing the benefit of leveraging deviation distribution over sending a fixed portion of tensors. Figure 9 shows the learned threshold for our method. The decreasing value of for indicates its focus on sending tensors with even smaller deviations during communication. As the model converges and becomes stable, the deviations decrease, resulting in generating a smaller threshold for a milder deviation distribution.
V-C2 Parameters used in MAML module
@@
The default structure of our MAML module is an MLP network with one hidden layer containing 100 neurons. To explore the impact of the number of neurons on performance, we keep the learning parameters consistent while varying the number of neurons in the hidden layer of the MAML module. In Figure 10, we see that using more neurons in the MAML module generally leads to improved results in terms of sacreBLEU score and tensor savings. The performance gain from using more neurons is intuitive since it provides additional degrees of freedom for learning optimization. However, it is important to note that using more neurons also incurs higher resource requirements during system construction.
V-C3 Meta evaluation for MAML module
@@
To examine the influence of the number of batches used to optimize our MAML module, Figure 11 shows the performance of our method with different numbers of batches used for meta-optimization. We see that increasing the number of samples used for MAML optimization generally results in improved translation quality and efficiency. Naturally, using more batches will increase the exact time spent on our MAML module. However, the time taken by clients to send parameters for aggregation may be more critical than this optimization time, as the optimization process is performed only once in each round.
V-C4 The use of pre-trained models
Tables VII and VIII show the sacreBLEU and COMET scores of each method without utilizing pre-trained knowledge for the M2M-100 model in each client. Specifically, each method was trained from scratch, utilizing randomly initialized weights. Although the results of do not exhibit complete superiority over as discussed in Section V-A, our MetaSend methods display higher average scores overall in most cases compared with baseline approaches. However, the overall results are relatively poor and unstable compared to the results shown in Tables I and II when pre-training is utilized. This discrepancy can be attributed to the fact that large neural MT systems typically demand a substantial amount of data and prior knowledge to perform effectively [49].